These are the news starting 01 Jan 2004 and ending December 2004.
Click here for the news of the previous years.
|19 Dec 04||The pump of the API ozone sensor ( ch.14) is defect. As a quick kludge, the API has been replaced by the old O341M sensor. Readings may be ca 20% too high...more on this asap|
|17 Dec 04||Maintenance by Envitec|
|10 Dec 04||Problems with the pump of the API sensor: very low O3 readings during the afternoon are a sign of a non-working pump.|
|06 Dec 04||Uploaded November 04 data.|
|24 Nov 04||
1. Maintenance and calibration work
2. The wet bulb sensor of the Heat-Stress device will not be rewetted for the next months, as we approach freezing temperatures and the concept of heat-stress becomes somewhat academic!
|01 Nov 04||Uploaded October 04 data.|
|13 Oct 04||Uploaded September 04 data.|
|06 Oct 04||The Hockey-Stick debate is heating up after von Storch's paper. We made a new special page to easy the access to these discussions; accessible through the main globalsceptics page or here!|
|02 Oct 04||Air-pressure sensor BS4 (channel 5) was 9.1 hPa low since many months (probably at least start of the year). An adjustement of 9.1 hPa has been made, taking as best reference sensors 2 Kestrel 4000 and a mercury barometer giving practically identical readings. See here for a graph.|
|26 Sep 04||New links on the "Today's Data" page to last 60 days sunspots (movie) and Antarctica ozone thickness (watch this for the ozone hole which will reach its greatest extend during the next 2 months (approx.))|
|25 Sep 04||Corrected CO2 connector. Some supplementary recalibration...|
|24 Sep 04||Maintenance work by Envitec. CO2 connector went loose...|
|11 Sep 04||Calibration and maintenance work on all sensors|
|09 Sep 04||Uploaded aug04.dat to archive. Resumed Dobson measurements with Microtops II.|
|15 Aug 04||Worlds Apart: new crystal-clear article on global warming by Prof. Philip Stott (thanks to Pierre Lutgen for forwarding)|
|14 Aug 04||Maintenance work on the sensors. Restarted radioactivity sensor.|
|01 Aug 04||Uploaded jul04.dat to archive|
|31 Jul 04||
Several problems during last 2 weeks:
1. radioactivity software again unreliable. Remedial must wait until September, alas.
2. the main postal modem driving the leased 2 Gbit/s line died (during heavy thunderstorm), so there were no live data available for about 2 days (22-23 July). New modem installed.
3. there is an obscure date problem with the ASCII files. All files have been restarted today, and the logger clock readjusted, so this is solved. Previous days graphs will be rebuild during the coming days.
4. the good news: no data have been lost!
5. look for the 20040723.mpg lightning movie for nice activity!
|12 Jul 04||Calibration of all gas sensors|
|10 Jul 04||
1. uploaded June data
2. the intermittent hanging of the "radioactivity computer" seem to be solved; the culprit was an extremely large log file.
|04 Jul 04||Uploaded jun04.dat file to archive.|
|02 Jul 04||CO2 sensor has become slightly instable (probably due to a IR-sensor/Peltier problem); the main readings still seem correct. A replacement CO2 sensor (from Edinburgh Sensors) has been mounted and connected to a second logger (these data are not displayed); this to not loose any data in case of the main sensor going down. The problem may be self-healing (or not...). Spikes are due to maintenance work.|
|25 Jun 04||Maintenance by Envitec on gas sensors.|
|11 Jun 04||Main Linux computer broke down again at 13:00 .. will be replaced tomorrow.|
|10 Jun 04||Official inaugurattion of the new location and equipments of meteoLCD|
|09 Jun 04||Main Linux computer up again.|
|08 Jun 04||
1. made a quick kludgy mounting to observe the
transit of Venus (look here!)
2. Main Linux computer break down at 18:30 UTC. No live data for the moment.
|07 Jun 04||The WBGT sensor placement was changed to a location facing East. The reason: the afternoon sunrays were partially reflected by an aluminium roofing and increased artificially globe, dry and wet bulb temperatures.|
|06 Jun 04||
1. The drybulb temperature of the WBGT sensor is
too high during high solar irradiances; replaced in graph of
WBGT by the usual air temperature sensor (the influence is
small, as the dry bulb temperature has a multiplier of only 0.1
in the calculation of the WBGT)
2. Uploaded may04.dat file
|07 May 2004||
Well...the movie making problems were not solved
after all. Most are related to the huge workload on the server
making the mpeg encoding (and are difficult to find). The
procedure was changed today: a movie is created every hour (it
remains invisible), and at midnight all these partial mpegs are
concatenated into the final accessible movie. The volume of the
movie could be heavily reduced in discarding every frame without
any activity, but this will completely upset smooth time-flow.
Every suggestion is welcome!
The latest version of the Windows Media Player 9 (build 9.00.00.2980) does not give the error described anymore... So let's hope that this will bring us into post-beta state!
|01 May 2004||Some (serious) teething problems with the lightning movie making... solved now! We should step into post-beta tomorrow! Microsoft Mediaplayer gives error when clicking on the file name, but works ok if movie is downloaded. Other players (Quicktime, ...) behave correctly!|
|30 Apr 2004||Started automatic making of lightning movies: each mpeg movie represents the activity for the day given as the file name. No sexy selectors for the moment, just pick your day! This is still early beta!|
|17 Apr 2004||A short report by Francis Massen on air temperature trends at meteoLCD (1998..2003)|
|03 Apr 2004||A computer shutdown blocked the transmission of live data from 03 April 08:00 to today 08:00. Ok now. Radiation sensor also back to work.|
|01 Apr 2004||The computer driving the radioactivity sensor is being upgraded...which will hopefully give a smoother operation...No radioactivity data for about 2 days!|
|21 Mar 2004||There was a longer power outage yesterday: no live-data available from 20/03 11:00 to 21/03 08:00. Back to normal today.|
|07 Mar 2004||New (link to) blockbuster report by Kyrill Kondratyev on the Globalsceptics page|
|29 Feb 2004||The data-logger can not handle the leap-year 29 Feb. Sorry, no data available today! The needed corrections (mainly clock resetting and fiddling with the data-files) will be applied tomorrow... no data will be lost.|
|16 Feb 2004||Changed webcam orientations so that pan dial is defined as follows: 0 points to true North, 90 is East, 180 is South, 270 (or -90) is West.|
|Uploaded January 2004 data file.|
|09 Feb 2004||Webcam now working as it should. To protect the privacy of the neighbours the zoom function is disabled and allowed only for some of the preset views on the sensors.|
|07 Feb 2004||
1. Steerable webcam running; Lisa software
from Seetec used to manage queues
and many parameters.
2. Radioactive sensor sometimes goes sleeping: the interface holds the last reading (see 7 days graphs)
3. New article of Francis Massen in the GlobalSceptics page: Êtes-vous Kyoto-crédule?.
|25 Jan 2004||
1. Updated readme
file with the new sensors; also pictures of all sensors in use
and links to the manufacturers.
2. The website now has a mirror: http://meteo01.lcd.lu; it is alpha state: data and graphs are correctly mirrored, but some pages are not!
|21 Jan 2004||
Logger halted to change
1. Ch12 = Radioactivity now holds dose-rate in nSv/h (and not mV)
2. Gruesome offsets of UVA_DT (Ch7) and UVB_DT (Ch8) corrected; take these two sensor readings with a pinch of salt; they are here as spare in the case of a major malfunction of the UV-Biometers
3. Corrected slight offset of solarrad sensor Ch23 (this is also a spare)
4. Restarted logging at 19:30. Be aware that 7days graphs may be some-how disrupted!
|08 Jan 2004||
BIG Event: Radiation Monitoring now online
(dose-rate of ambient gamma radiation). This was a technological
challenge, as the sensor is pure RS232 based and we need a
voltage proportional to the dose-rate. Claude
Baumann, a top expert in programmable microcontrollers,
built and programmed the needed device... which worked from 1st
second on! See here
Claude's robotic site.
For the moment we make calibration tests with radioactive samples: the data shown DO NOT represent ambient radioactivity for the moment !
|03 Jan 2004||First preliminary comparison between the KMI Brewer and the meteoLCD total ozone column data. The KMI data have been "unscanned" from their today's graph (full year 2003).|
|02 Jan 2004||The usual New Year problem with the 7 days graph (the logger data do not contain the year information, so the automatically graphing by Gnuplot gets confused). Restarted 7days graphs at 01Jan2004|